# Random Walk on Pixel Manifolds for Anomaly Segmentation of Complex Driving Scenes

### v1

@@ Abstract
In anomaly segmentation for complex driving scenes, state-of-the-art approaches utilize anomaly scoring functions to calculate anomaly scores. For these functions, accurately predicting the logits of inlier classes for each pixel is crucial for precisely inferring the anomaly score. However, in real-world driving scenarios, the diversity of scenes often results in distorted manifolds of pixel embeddings in embedding space. This effect is not conducive to directly using the pixel embeddings for the logit prediction during inference, a concern overlooked by existing methods. To address this problem, we propose a novel method called Random Walk on Pixel Manifolds (RWPM). RWPM utilizes random walks to reveal the intrinsic relationships among pixels to refine the pixel embeddings. The refined pixel embeddings alleviate the distortion of manifolds, improving the accuracy of anomaly scores. Our extensive experiments show that RWPM consistently improve the performance of the existing anomaly segmentation methods and achieve the best results. Code: https://github.com/ZelongZeng/RWPM.
@@ 1 Introduction
Figure 1: A toy example about manifolds of pixel embeddings. The dashed circles indicate regions with high predicted logit to the corresponding inlier class. For the pixel in these regions, its anomaly score is small. The prototype is the vector of network weights of the classifier for the corresponding inlier class. (i) An ideal manifold structure. All inlier pixels are within the corresponding high logit region, whereas all outlier pixels are outside the area. The anomaly score derived from the anomaly scoring function effectively discriminates between them. (ii) The manifolds are affected by the diversity of data, causing some inlier pixels to deviate from the corresponding high logit region, while some outlier pixels approach the regions. This results in false positives/negatives in the anomaly scores. However, pixels of the same class are still in the same manifold, indicating that the manifold structure can be utilized to reveal the intrinsic relationships between pixels. (iii) Our RWPM utilizes random walks to capture the manifold structure to diffuse and update pixel embeddings. The embeddings within the same manifold tend to become more similar after the updating. Fig 3 and Fig 4 respectively present the visualization of embedding distributions and the qualitative results of real examples, consistent with the description of this toy example.
@@
The anomaly scoring functions in OE methods infer anomaly scores based on the inlier class logit predictions from a semantic segmentation model. Specifically, the logic of these functions can be represented as follows: when the predicted logits of a pixel embedding are small across all inlier classes, indicating that the pixel is dissimilar to all inlier classes and is an outlier with high probability, the function yields a higher anomaly score. Therefore, accurately measuring the logits of each pixel is crucial for precisely evaluating the anomaly score. However, in practical open-world driving scenarios, various environmental conditions (such as lighting, road materials, locations, etc.) and a diverse range of unknown class objects are present. During the inference phase, the diversity of driving scenes often result in distorted manifolds of pixel embedding in the embedding space, causing pixel embeddings to deviate from the ideal distribution in the embedding space. This phenomenon has been validated in previous works of metric learning [3, 25, 41, 42]. Since the logits are calculated from the inner product (i.e., measuring cosine distance) of the pixel embeddings and the vectors of network weights of classifiers, the distorted manifolds is not conducive to directly using the pixel embeddings for the logit prediction. However, current anomaly segmentation methods ignore this issue and directly use the pixel embeddings for the logit prediction in the inference phase, so the results often fail to accurately reflect the anomaly score for the corresponding pixels, causing high false positive/negative rate (see Fig. 1(i) and Fig. 1(ii)).
@@
In this paper, we propose a novel method, called Random Walk on Pixel Manifolds (RWPM), to alleviate the effect of data diversity on the manifolds. RWPM utilizes random walks to capture the manifold structure of the pixel embeddings to diffuse and update the embedding of each pixel. Since the structure of data manifolds implicitly contains the intrinsic relationship among data points [44, 3, 24, 41], diffusing and updating pixel embeddings on the pixel manifolds gradually leads to increased similarity among the embeddings on the same manifold. The updated pixel embedding can be seen as a weighted ensemble of other pixel embeddings, with the weights determined by their intrinsic relationship. In the embedding space, this manifests as the pixel embeddings on the same manifold forming a more compact cluster, thereby mitigating the effect of data diversity on the pixel manifolds, which facilitates the logit prediction (see Fig. 1(iii)). Furthermore, since the typically large number of pixels in an image, utilizing all pixels for the random walks results in significant memory and computational overhead, which damages the deployment and real-time execution of our RWPM. To address this issue, we introduce the Partial Random Walk strategy for RWPM, reducing memory consumption and enhancing the operational efficiency of RWPM. Note that, RWPM is proposed to use in the inference phase, thus no any additional training is required. Additionally, RWPM can be directly integrated into existing anomaly detection framework without the need for any extra modification to the network structure. Extensive experiments in various anomaly segmentation benchmarks of road scenes, namely Fishyscapes [5], Road Anomaly [31], and SMIYC [6], demonstrate that RWPM consistently improves the performance of existing methods, achieving SOTA results. To summarise, our contributions are the following:
@@
  • We propose the Partial Random Walk strategy for RWPM to reduce memory consumption and improve operational efficiency.
@@ 2 Related Work
Anomaly Segmentation. Anomaly segmentation aims to detect the outlier class object in some specific scenes, such as complex driving scenes [13, 31, 6]. Existing anomaly segmentation can be broadly divided into three categories: (a) Uncertainty-based methods, (b) Reconstruction-based methods and (c) Outlier Exposure methods. Uncertainty-based methods assume that outlier samples result in low-confidence predictions. Based on this assumption, they focus on measuring the pixel-wise anomaly score by estimating the prediction uncertainty through various ways, such as ensembles [27], Bayesian estimation [16], maximum softmax probability [21, 29], and logit [20, 26]. Since segmentation models are trained on close-set, they may misclassify unseen classes with high confidence, leading uncertainty-based methods fail to detect anomalies. Reconstruction-based methods detect anomaly pixel by comparing the discrepancy between the original image and its reconstructed image generated based on the semantic segmentation predictions [14, 31, 36, 15]. These methods are challenged by the accuracy of segmentation prediction, and by the difficulties of training. Recent methods [7, 38, 12, 37, 35, 19] introduce the Outlier Exposure (OE) strategy [22] that utilizes auxiliary outlier data and anomaly scoring functions to learn models. The auxiliary outlier data is the image from other datasets (e.g., COCO [30]), or cutting the outlier objects from other datasets and pasting them into the inlier scenes, and the anomaly scoring functions are used to calculate the anomaly score for each pixel. Compared with previous works, OE strategy can reach better performance. More recently, some OE methods [37, 35, 19] adopt mask-based segmentation networks [11, 10], instead of pixel-based segmentation networks [8, 9], as their backbones. Based on the powerful capability of mask-based networks in segmentation tasks, these methods significantly reduces false detection rates, achieving SOTA results. In this paper, we only concentrate on the OE methods. The anomaly scoring functions employed in OE methods are based on the predicted logits of inlier classes. However, during the inference phase, the diversity of driving scenes results in distorted manifolds of pixel embeddings. As a result, the scoring functions may fail to use the embeddings to accurately calculate logits, thereby affects the accuracy of measuring anomaly scores. Current OE methods ignore this problem. Our proposed RWPM addresses this problem by utilizing random walks on pixel manifolds to reveal the intrinsic relationships among pixels to refine the pixel embeddings. To the best of our knowledge, RWPM is the first work to propose and resolve this problem.
@@ 3.1 Preliminaries
For anomaly scoring functions, their logic can be represented as follows: when the predicted logits of a pixel embedding are small across all inlier classes, indicating that the pixel is dissimilar to all inlier classes and is likely an outlier, the function generate a higher anomaly score. For example, let’s define $\boldsymbol{l}_{h,w}\left(k\right)$ is the logit of a pixel $\boldsymbol{x}_{h,w}$ for class $k$. The scoring function in PEBAL [38] is $E_{h,w}=-\log\sum_{k\in\left\{1\cdots K\right\}}\exp\left(\boldsymbol{l}_{h,w}\left(k\right)\right)$. Obviously, when the anomaly score $E_{h,w}$ is large, it indicates that the logits of the pixel across all inlier classes are small, i.e., the pixel $\boldsymbol{x}_{h,w}$ is dissimilar from all inlier classes (refer to Section 6.2 of the supplementary section for the analysis of other anomaly scoring functions).
@@ 3.2 Graph Construction
where $\tau$ is a temperature hyper-parameter, we set its value less than $1.0$ to remove the effect of non-neighbor points. Compared to typical search algorithms, Eq. 2 can impose local constraints more efficiently and effectively (see Section 6.3 of the supplementary), and keep the scale of embedding in subsequent random walks.
@@ 3.3 Random Walk Process
Now we have constructed the graph $\mathbf{S}$, we can start the random walk. Different from previous mining on data manifolds works utilizing random walks to propagate the states of known samples to other unknown samples, we use a random walk to propagate the embedding between pixels on manifolds.
@@ 3.4 Partial Random Walk
Next, for each sub-map $\boldsymbol{p}^{(i)}$, we perform random walks with the limited iteration strategy in their corresponding sub-graphs $\boldsymbol{S}_{\boldsymbol{p}^{(i)}}$. That is, we use Eq. 3 for a small number of iterations instead of using the closed-form Eq. 4. We found that a few iterations by using Eq. 3 are enough to yield high quality results. The time complexity of Eq. 3 with $T$ iterations for each sub-map is $O(Td(\frac{HW}{n^{2}})^{2})$, while the time complexity of Eq. 4 is $O((\frac{HW}{n^{2}})^{3})+O(d(\frac{HW}{n^{2}})^{2})$. Since $Td$ is much smaller than $\frac{HW}{n^{2}}$, the limited iteration strategy can significantly improve the efficiency of RWPM. After the random walks, we obtain the refined pixel embeddings sub-map $\mathbf{m}^{T}_{\boldsymbol{p}^{(i)}}$ for each sub-map $\boldsymbol{p}^{(i)}$. Finally, we concatenate the refined sub-maps based on their positions to obtain the final refined pixel embedding map. Additionally, When $n$ is large (observed experimentally as $n>2$), due to spatial constraints, the baseline values of anomaly scores outputted on each sub-map may vary. Hence, a calibration is performed. For adjacent refined sub-maps $\boldsymbol{p}^{(i)^{\prime}}$ and $\boldsymbol{p}^{(j)^{\prime}}$, we compute the average anomaly scores $I$ and $J$ of all pixels on their adjacent edges, respectively. Subsequently, using $\boldsymbol{p}^{(i)^{\prime}}$ as the reference, we calculate the ratio $\frac{I}{J}$, and subsequently multiply all anomaly scores on sub-map $\boldsymbol{p}^{(j)^{\prime}}$ by this ratio to achieve the calibration.
@@ 4 Experiment
Evaluation Metrics: Following the previous works, we evaluate our method on the standard metrics for anomaly segmentation. On the Fishyscapes Lost&Found and Road Anomaly, we adopt the the area under receiver operating characteristics (AUROC), average precision (AP), and the false positive rate at a true positive rate of 95% (FPR95). On the SMIYC, in addition to reporting per-pixel metrics such as AP, AUC, and FPR95, we also calculate results for component-level metrics [6] such as the component-wise intersection over union (sIoU), positive predictive value (PPV), and mean F1.
@@ 4.1 Effectiveness of RWPM
First, to validate that RWPM can improve the performance of existing anomaly segmentation methods, we integrate our approach with four recent representative methods, namely PEBAL [38], Balanced Energy [12], Mask2Anomaly [37] and RbA [35]. PEBAL and Balanced Energy employ pixel-based networks as their backbones, whereas Mask2Anomaly and RbA utilize mask-based networks as their backbones. We evaluate the results on both the Fishyscapes Lost&Found validation set and Road Anomaly test set. The experimental results are reported in Table 1. These results demonstrate that our proposed RWPM consistently and significantly enhances the performance of existing anomaly segmentation methods, across different backbone architectures.
Second, to further demonstrate the effectiveness of our RWPM, we compare it with other state-of-the-art (SOTA) methods. Specifically, we adopt the combination of RbA with RWPM as our instance approach for the comparison. In addition to the Fishyscapes and Road Anomaly datasets, we also evaluate our approach on a more challenging dataset, i.e., SMIYC online test set (Anomlay track and Obstacle track). SMIYC is characterized by its high domain shift and diversity of objects. Table 2 presents the results. The results indicate that our method can consistently achieve high performance across various datasets, while the previous SOTA methods fail to do so. We also compare all approaches by computing the average results of AP and FPR95. The results shows that our approach outperforms the previous methods and achieve the best results.
@@
PEBAL [38] Pixel-base 98.96 58.81 4.77 87.63 45.10 44.58
@@
Balanced Energy [12] Pixel-base 99.03 67.07 2.93 88.31 41.48 41.46
@@
Mask2Anomaly [^1] We carefully reproduce the experiments by using the official code. [37] Mask-base 95.41 69.44 9.22 96.54 80.04 13.95
Mask2Anomaly + RWPM (Ours) Mask-base 95.92 69.54 8.13 97.44 80.09 7.45
RbA [35] Mask-base 98.62 70.81 6.30 97.99 85.42 6.92
@@
RbA [35] (ICCV’23) Mask-based 70.81 6.30 85.42 6.92 90.9 11.6 91.8 0.5 84.73 6.33
RbA + RWPM (Ours) Mask-based 71.16 6.12 87.34 5.27 92.20 10.15 93.30 0.28 86.00 5.46
@@
RbA + RWPM (Ours) Mask-based 57.00 61.25 58.44 59.47 72.51 69.85

### v2

@@ Abstract
In anomaly segmentation for complex driving scenes, state-of-the-art approaches utilize anomaly scoring functions to calculate anomaly scores. For these functions, accurately predicting the logits of inlier classes for each pixel is crucial for precisely inferring the anomaly score. However, in real-world driving scenarios, the diversity of scenes often results in distorted manifolds of pixel embeddings in the space. This effect is not conducive to directly using the pixel embeddings for the logit prediction during inference, a concern overlooked by existing methods. To address this problem, we propose a novel method called Random Walk on Pixel Manifolds (RWPM). RWPM utilizes random walks to reveal the intrinsic relationships among pixels to refine the pixel embeddings. The refined pixel embeddings alleviate the distortion of manifolds, improving the accuracy of anomaly scores. Our extensive experiments show that RWPM consistently improve the performance of the existing anomaly segmentation methods and achieve the best results. Code is available at: https://github.com/ZelongZeng/RWPM.
Abstract
The following items are included in the supplementary material.
    The Architecture of Pixel-based and Mask-based Networks.
    Anomaly Scoring Function Details.
    Additional Results and Implementation Details.
    Different Methods for Constructing Locally Constrained Graph.
    Analysis for RWPM.
    Some Reviewers’ Comments and Our Explanations.
@@ 1 Introduction
Figure 1: A toy example about manifolds of pixel embeddings. The dashed circles indicate regions with high predicted logit to the corresponding inlier class. For the pixel in these regions, its anomaly score is small. The prototype is the vector of network weights of the classifier for the corresponding inlier class. (i) An ideal manifold structure. All inlier pixels are within the corresponding high logit region, whereas all outlier pixels are outside the areas. The anomaly score derived from the anomaly scoring function effectively discriminates between them. (ii) The manifolds are affected by the diversity of data, causing some inlier pixels to deviate from the corresponding high logit region, while some outlier pixels approach the regions. This results in false positives/negatives in the anomaly scores. However, pixels of the same class are still in the same manifold, indicating that the manifold structure can be utilized to reveal the intrinsic relationships between pixels. (iii) Our RWPM utilizes random walks to capture the manifold structure to diffuse and update pixel embeddings. The embeddings within the same manifold tend to become more similar after the updating. Fig 3 and Fig 4 respectively present the visualization of embedding distributions and the qualitative results of real examples, consistent with the description of this toy example.
@@
The anomaly scoring functions in OE methods infer anomaly scores based on the inlier class logit predictions from a semantic segmentation model. Specifically, the logic of these functions can be represented as follows: when the predicted logits of a pixel embedding are small across all inlier classes, indicating that the pixel is dissimilar to all inlier classes and is an outlier with high probability, the function yields a higher anomaly score. Therefore, accurately measuring the logits of each pixel is crucial for precisely evaluating the anomaly score. However, in practical open-world driving scenarios, various environmental conditions (such as lighting, road materials, etc.) and a diverse range of unknown class objects are present. During the inference phase, the diversity of driving scenes often results in distorted manifolds of pixel embedding in the embedding space, causing pixel embeddings to deviate from the ideal distribution. This phenomenon has been validated in previous works of metric learning [3, 25, 41, 42]. Since the logits are calculated from the inner product (i.e., cosine distance) of the pixel embeddings and the vectors of classifiers’ network weights, the distorted manifolds are not conducive to directly using the pixel embeddings for the logit prediction. However, current anomaly segmentation methods ignore this issue and directly use the pixel embeddings for the logit prediction in the inference phase, so the results often fail to accurately reflect the anomaly score for the corresponding pixels, causing high false positive/negative rate (see Fig. 1(i) and Fig. 1(ii)).
@@
In this paper, we propose a novel method, called Random Walk on Pixel Manifolds (RWPM), to alleviate the effect of data diversity on the manifolds. RWPM utilizes random walks to capture the manifold structure of the pixel embeddings to diffuse and update the embedding of each pixel. Since the structure of data manifolds implicitly contains the intrinsic relationship among data points (see Fig. 1(ii)) [44, 3, 24, 41], using random walks to measure the similarity between pixels on manifolds (i.e., distance on manifolds rather than Euclidean distance) better reflects the intrinsic relationships between pixels. Consequently, diffusing and updating pixel embeddings based on the similarity on manifolds results in high similarity among pixels on the same manifold and low similarity among pixels on different manifolds. In the embedding space, this manifests as pixel embeddings on the same manifold forming more compact clusters, thereby mitigating the distortion effect of data diversity on the manifolds, which facilitates the logit prediction (see Fig. 1(iii)). Furthermore, since the typically large number of pixels in an image, utilizing all pixels for the random walks results in significant memory and computational overhead, which damages the deployment and real-time execution of our RWPM. To address this issue, we introduce the Partial Random Walk strategy for RWPM, reducing memory consumption and enhancing the operational efficiency of RWPM. Note that, RWPM is proposed to use in the inference phase, thus no any additional training is required. Additionally, RWPM can be directly integrated into existing anomaly segmentation framework without the need for any extra modification to the network structure. Extensive experiments in various anomaly segmentation benchmarks of road scenes, namely Fishyscapes [5], Road Anomaly [31], and SMIYC [6], demonstrate that RWPM consistently improves the performance of existing methods, achieving SOTA results. To summarise, our contributions are the following:
@@
  • We propose the Partial Random Walk strategy for RWPM to reduce the memory consumption and improve the operational efficiency.
@@ 2 Related Work
Anomaly Segmentation. Anomaly segmentation aims to detect the outlier class object in some specific scenes, such as complex driving scenes [13, 31, 6]. Existing anomaly segmentation can be broadly divided into three categories: (a) Uncertainty-based methods, (b) Reconstruction-based methods and (c) Outlier Exposure methods. Uncertainty-based methods assume that outlier samples result in low-confidence predictions. Based on this assumption, they focus on measuring the pixel-wise anomaly score by estimating the prediction uncertainty through various ways, such as ensembles [27], Bayesian estimation [16], maximum softmax probability [21, 29], and logit [20, 26]. Since segmentation models are trained on close-set, they may misclassify unseen classes with high confidence, leading uncertainty-based methods fail to detect anomalies. Reconstruction-based methods detect anomaly pixel by comparing the discrepancy between the original image and its reconstructed image generated based on the semantic segmentation predictions [14, 31, 36, 15]. These methods are challenged by the accuracy of segmentation prediction and the difficulties of training. Recent methods [7, 38, 12, 37, 35, 19] introduce the Outlier Exposure (OE) strategy [22] that utilizes auxiliary outlier data and anomaly scoring functions to learn models. The auxiliary outlier data is the image from other datasets (e.g., COCO [30]), or cutting the outlier objects from other datasets and pasting them into the inlier scenes, and the anomaly scoring functions are used to calculate the anomaly score for each pixel. Compared with previous works, OE strategy can reach better performance. More recently, some OE methods [37, 35, 19] adopt mask-based segmentation networks [11, 10], instead of pixel-based segmentation networks [8, 9], as their backbones. Based on the powerful capability of mask-based networks in segmentation tasks, these methods significantly reduces false detection rates, achieving SOTA results. In this paper, we only concentrate on the OE methods. The anomaly scoring functions employed in OE methods are based on the predicted logits of inlier classes. However, during the inference phase, the diversity of driving scenes results in distorted manifolds of pixel embeddings. As a result, the scoring functions may fail to use the embeddings to accurately calculate logits, thereby affects the accuracy of measuring anomaly scores. Current OE methods ignore this problem. Our proposed RWPM addresses this problem by utilizing random walks on pixel manifolds to reveal the intrinsic relationships among pixels and refine the pixel embeddings. To the best of our knowledge, RWPM is the first work to propose and resolve this problem.
@@ 3.1 Preliminaries
For anomaly scoring functions, their logic can be represented as follows: when the predicted logits of a pixel embedding are small across all inlier classes, indicating that the pixel is dissimilar to all inlier classes and is likely an outlier, the function generates a higher anomaly score. For example, let’s define $\boldsymbol{l}_{h,w}\left(k\right)$ is the logit of a pixel $\boldsymbol{x}_{h,w}$ for class $k$. The scoring function in PEBAL [38] is $E_{h,w}=-\log\sum_{k\in\left\{1\cdots K\right\}}\exp\left(\boldsymbol{l}_{h,w}\left(k\right)\right)$. Obviously, when the anomaly score $E_{h,w}$ is large, it indicates that the logits of the pixel across all inlier classes are small, i.e., the pixel $\boldsymbol{x}_{h,w}$ is dissimilar from all inlier classes (see Section S.2 of the supplementary material for the analysis of other anomaly scoring functions).
@@ 3.2 Graph Construction
where $\tau$ is a temperature hyper-parameter, we set its value less than $1.0$ to remove the effect of non-neighbor points. Compared to typical search algorithms, Eq. 2 can impose local constraints more efficiently and effectively (see Section S.4 of the supplementary material), and keep the scale of embedding in subsequent random walks.
@@ 3.3 Random Walk Process
Now we have constructed the graph $\mathbf{S}$, we can start the random walks. Different from previous mining on data manifolds works utilizing random walks to propagate the states of known samples to other unknown samples, we use a random walk to propagate the embedding between pixels on manifolds.
@@ 3.4 Partial Random Walk
Next, for each sub-map $\boldsymbol{p}^{(i)}$, we perform random walks with the limited iteration strategy in their corresponding sub-graphs $\boldsymbol{S}_{\boldsymbol{p}^{(i)}}$. That is, we use Eq. 3 for a small number of iterations instead of using the closed-form Eq. 4. We found that a few iterations by using Eq. 3 are enough to yield high quality results. The time complexity of Eq. 3 with $T$ iterations for each sub-map is $O(Td(\frac{HW}{n^{2}})^{2})$, while the time complexity of Eq. 4 is $O((\frac{HW}{n^{2}})^{3})+O(d(\frac{HW}{n^{2}})^{2})$. Since $Td$ is much smaller than $\frac{HW}{n^{2}}$, the limited iteration strategy can significantly improve the efficiency of RWPM. After the random walks, we obtain the refined pixel embeddings sub-map $\mathbf{m}^{T}_{\boldsymbol{p}^{(i)}}$ for each sub-map $\boldsymbol{p}^{(i)}$. Finally, we concatenate the refined sub-maps based on their positions to obtain the final refined pixel embedding map. Additionally, When $n$ is large (observed experimentally as $n>2$), due to spatial constraints, the baseline values of anomaly scores outputted on each sub-map may vary. Hence, a calibration is performed. For adjacent refined sub-maps $\boldsymbol{p}^{(i)^{\prime}}$ and $\boldsymbol{p}^{(j)^{\prime}}$, we compute the average anomaly scores $I$ and $J$ of all pixels on their adjacent edges, respectively. Subsequently, using $\boldsymbol{p}^{(i)^{\prime}}$ as the reference, we calculate the ratio $\frac{I}{J}$, and then multiply all anomaly scores on sub-map $\boldsymbol{p}^{(j)^{\prime}}$ by this ratio to achieve the calibration.
@@ 4 Experiment
Evaluation Metrics: Following the previous works, we evaluate our method on the standard metrics for anomaly segmentation. On the Fishyscapes Lost&Found and Road Anomaly, we adopt the the area under receiver operating characteristics (AUROC), average precision (AP), and the false positive rate at a true positive rate of 95% (FPR95). On the SMIYC, in addition to reporting per-pixel metrics such as AP and FPR95, we also calculate results for component-level metrics [6] such as the component-wise intersection over union (sIoU), positive predictive value (PPV), and mean F1.
@@ 4.1 Effectiveness of RWPM
First, to validate that RWPM can improve the performance of existing anomaly segmentation methods, we integrate our approach with four recent representative methods, namely PEBAL [38], Balanced Energy [12], Mask2Anomaly [37] and RbA [35]. PEBAL and Balanced Energy employ pixel-based networks as their backbones, whereas Mask2Anomaly and RbA utilize mask-based networks as their backbones. We evaluate the results on both the Fishyscapes Lost&Found validation set and Road Anomaly test set. The experimental results are reported in Table 1. These results demonstrate that our proposed RWPM consistently and significantly enhances the performance of existing anomaly segmentation methods, across different backbone architectures. Additional results and parameters details are shown in Section S.3 of the supplementary material.
@@
PEBAL [38] Pixel-based 98.96 58.81 4.77 87.63 45.10 44.58
@@
Balanced Energy [12] Pixel-based 99.03 67.07 2.93 88.31 41.48 41.46
@@
Mask2Anomaly [^2] We carefully reproduce the experiments by using the official code. [37] Mask-based 95.47 65.27 7.79 96.54 80.04 13.95
Mask2Anomaly + RWPM (Ours) Mask-base 95.48 65.38 7.33 97.44 80.09 7.45
RbA [35] Mask-based 98.62 70.81 6.30 97.99 85.42 6.92
@@
RbA [35] (ICCV’23) Mask-based 70.81 6.30 85.42 6.92 90.86 11.59 91.85 0.46 84.73 6.33
RbA + RWPM (Ours) Mask-based 71.16 6.12 87.34 5.27 92.00 10.15 93.30 0.28 86.00 5.46
Second, to further demonstrate the effectiveness of our RWPM, we compare it with other state-of-the-art (SOTA) methods. Specifically, we adopt the combination of RbA with RWPM as our instance approach for the comparison. In addition to the Fishyscapes and Road Anomaly datasets, we also evaluate our approach on a more challenging dataset, i.e., SMIYC online test set (Anomaly track and Obstacle track). SMIYC is characterized by its high domain shift and diversity of objects. Table 2 presents the results. The results indicate that our method can consistently achieve high performance across various datasets, while the previous SOTA methods fail to do so. We also compare all approaches by computing the average results of AP and FPR95. The results shows that our approach outperforms the previous methods and achieve the best results.
@@
RbA + RWPM (Ours) Mask-based 57.00 61.25 58.44 58.89 72.51 69.85